OpenStudy (anonymous):

Find the constant difference for a hyperbola with foci at F1 (0, -10) and F2 (0, 10) and the point on the hyperbola (6, 7.5)

OpenStudy (anonymous):

So I plugged in the numbers for the formula |√x1-x3)^2+(y1-y3)^2 -√(x2-x3)^2+(y2-y3)^2| Which would be: |√0-6)^2+(-10-7.5)^2 -√(0-6)^2+(10-7.5)^2|

OpenStudy (jdoe0001):

that's correct

OpenStudy (anonymous):

Ok so when I simplified the problem it looks like this: √-6^2+-17.5^2 - √-6^2+2.5^2

OpenStudy (jdoe0001):

\(\bf (0-6)^2\implies (-6)^2\implies -6\cdot -6\implies +36\)

OpenStudy (anonymous):

So then I got √36+306.25 - √36+6.25

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Which would end up as -16.439-42.25 It doesn't look right I don't think it is the answer

OpenStudy (anonymous):

-58.689 ?

OpenStudy (jdoe0001):

well \(\bf \sqrt{(0-6)^2+(-10-7.5)^2}-\sqrt{(0-6)^2+(10-7.5)^2} \\ \quad \\ \sqrt{(-6)^2+(-17.5)^2}-\sqrt{(-6)^2+(-2.5)^2} \\ \quad \\ \sqrt{36+306.25}-\sqrt{36+6.25} \\ \quad \\ \sqrt{342.25}-\sqrt{42.25}\)

OpenStudy (anonymous):

18.5-6.5=12

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Thank you! I appreciate your help

OpenStudy (jdoe0001):

yw